Systems Staff Meeting Notes 5/22/00

Systems Staff Meeting Notes
May 22, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
  1. The Miami Heat lost their final game of the season over the weekend on a controversial call.
  2. Systems shirts have arrived! They were a big hit at the MMM last week.
  3. The A&P DLC Programmer position closes this Friday.
  4. Stacy Wall is leaving as of June 1. She's taking a position at the YMCA camp in Keystone Heights.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We've received 2 laptops that were ordered for West Reference and the East Conference Room. We need to make sure that they are equipped with Xircom modem cards. If they did not come with the cards, then we need to order them.
  2. Some orders have been held up in Accounting. They are checking on money availability.
  3. A $7000 Incoming/Outgoing Mail Scanner has been ordered for the mail server. It currently uses NetShield, but this new scanner will provide better mail filtering ability and will also protect internally. For example, WAZOO keeps getting re-introduced. This scanner should protect against such occurrences.
  4. Orders for Systems furniture have been placed: a 12-seat replacement for the conference table, Will's ergonomic U-shaped desk, Debra & the students' cubicle area. Russ told Debra that Bill would need to sign off on all these orders. Bill was unaware of this fact.
  1. Liaisons have been scurrying to do the virus/Office updates.
  2. Everett notes that Dot Hope's machine would not give her the Windows Critical Update information after running the office patch. Bill suggests checking to see what security level her machine is set at.
  3. Some machines have been asking for Disk 2 of O2K. Maybe on new machines where there was a factory install of Office as opposed to the download from the server. Need to find out what type of machine is experiencing this problem.
  1. The Xerox IPX problem was solved by rebooting the server (twice).
  2. We have not yet received the floor plans detailing the DLC move. DLC is reportedly moving to the 2nd floor and some of Resource Services will move up to the 3rd floor.
  3. West wiring - new dated of completion projected to be Thursday, June 1.
  4. Novell conversion - waiting on software.
  5. Wireless - waiting on Brad & Digital.
  6. TerraByte server - waiting for Bill to install Apache.
  7. Journalism expansion is just beginning. They are negotiating for more work stations. They may end up sharing domains. Xerox is not in place yet. Gary says they are still evaluating the situation.
Software Projects Status
  1. Will & Winston attempted some NT hot fixes this morning. Erich got caught with some open files - Winston had to restore them. NT3 is now at 6A.
  2. End of year accounting jobs to run next month.
  3. Ying has been working on the Systems wiring print out design. She will demo it for the hardware group this afternoon.
  4. Preservation's Q&A database will be demoed tomorrow. They are requesting two demos before the data transfer.
  5. Dot Hope - OPS tracking report requests - she is pretty happy with it. They have had about 50% compliance with their people doing the data entry required. They have considered having the students enter the data on their own for a more immediate response. Originally they did not want students to enter their own information - Bill would like to mull over this with Winston.
  6. The printer device database for Special Collections is waiting on Jeffrey to complete the data entry.
  7. Student Patron Updates are complete for Summer A/C.
  1. Hilton is out today.
  2. Debra out May 26.
  3. Memorial Day - Monday, May 29.
  4. Winston out June 19-23.
  5. Bobbie out July 10-21.
  6. Will out July 24-28.

Back to the Systems Homepage

Last updated June 5, 2000
by Debra Harris